OpenStudy (supersquidsisters):

For an angle θ with the point (15, –8) on its terminating side, what is the value of cosine? negative 8 over 18 15 over 17 –8 15

OpenStudy (lord_box):

Do you know the formula for cosine?

OpenStudy (supersquidsisters):

@Lord_Box cosine=adjacent/hypotenuse , right?

OpenStudy (lord_box):

Right. So if you have an angle at (15, -8) then it must go in which direction (right/left 15) and (up/down 8)

OpenStudy (supersquidsisters):

@Lord_Box um i still kinda dont understand how to get adjacent and hypotenuse,,, sorry

OpenStudy (lord_box):



OpenStudy (supersquidsisters):

@Lord_Box oh, so it would be cos = 15/-8 ??

OpenStudy (lord_box):

cos is adjacent/hypotenuse, which side is adjacent to the angle?

OpenStudy (supersquidsisters):

@Lord_Box so its -8/15 ????

OpenStudy (lord_box):

Again, which side is adjacent to the angle?

OpenStudy (supersquidsisters):

@Lord_Box uum the 15 side thats on the x line??

OpenStudy (lord_box):

Yes. That's adjacent. Now, how could we find the hypotenuse of a triangle with sides -8 and 15?

OpenStudy (supersquidsisters):

@Lord_Box umm no idea im so sorry :c

OpenStudy (lord_box):

Are you familiar with the Pythagorean theorem?

OpenStudy (supersquidsisters):

@Lord_Box yeah

OpenStudy (lord_box):

a^2 + b^2 = c^2, where c is the length of the hypotenuse. Plug in -8 and 15 for a and b.

OpenStudy (supersquidsisters):

@Lord_Box -8^2 + 15^2 = c^2 -64 + 225 = c^2 161 = c^2 ????

OpenStudy (lord_box):

make sure you're squaring (-8) and not - 8

OpenStudy (lord_box):

\[(-8)^2 = 64, - (8^2)=-64\]

OpenStudy (supersquidsisters):

@Lord_Box oh ok sorry

OpenStudy (lord_box):

Try the Pythagorean theorem again

OpenStudy (supersquidsisters):

with (-8)^2 now? @Lord_Box

OpenStudy (lord_box):

yes

OpenStudy (supersquidsisters):

@Lord_Box (-8)^2 + 15^2 = c^2 64 + 225 = c^2

OpenStudy (supersquidsisters):

289 = c^2 is this right so far?

OpenStudy (lord_box):

yes

OpenStudy (supersquidsisters):

ok so what now?

OpenStudy (supersquidsisters):

@Lord_Box ??

OpenStudy (lord_box):

What did you get for c?